# An All-Reduce Compatible Top-$K$ Compressor for Communication-Efficient Distributed Learning

### v2

@@ I Introduction
The rise of large-scale machine learning has established distributed training as a fundamental paradigm in modern AI systems. In centralized cloud environments, massive datasets are partitioned and sharded across NN nodes (e.g., GPUs) within a data center to enable data-parallel training [1] [2]. In contrast, federated learning keeps data on NN client devices (e.g., smartphones) for privacy and regulatory compliance, while a coordinating server aggregates model updates instead of raw data [3, 4, 5]. Both paradigms give rise to the following distributed stochastic optimization problem:
@@
Here, NN denotes the number of computing nodes, $\bm{\xi}$ represents a random data sample drawn from the local distribution $\mathcal{D}_{i}$, and $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the local loss function at node $i$, which may be non-convex. While each node possesses its local data and loss function, they collaborate to minimize the global loss function $f(\bm{x})$ defined in problem (1).
In such distributed settings, each node $i$ can only access its local data and calculate local gradient $\nabla F_{i}(\bm{x};\bm{\xi})$ of its loss function during optimization. However, communication is required to obtain information from other nodes, which frequently becomes the primary bottleneck and dominates end-to-end training time. The associated cost scales with both the model dimension $d$ and the number of nodes NN, rendering it prohibitive for large-scale learning tasks. Consequently, minimizing the volume of communicated bits per round is a critical problem for efficient and scalable distributed learning.
@@ I-B Main Results
Table I compares Rand-$K$, Top-$K$, and ARC-Top-$K$. The results show that ARC-Top-$K$ enables communication-efficient algorithms to achieve the same theoretical convergence rate and communication overhead as Rand-$K$, while matching the empirical performance of Top-$K$. All reported convergence rates are obtained in combination with EF21M [8], and the transient complexity reflects the number of iterations required for the asymptotic rate term to dominate. In the table, NN is the number of nodes, $T$ is the number of iterations, $\sigma$ is the standard deviation of the stochastic gradient noise. Tensors in all algorithm are compressed in 2-dimensional view of $m\times n$ while $K$ represents the number of selected rows in each tensor.
@@ II Related Works
Distributed Learning. Distributed optimization underpins large-scale training in both data-center clusters and cross-device networks [1] [2] [9, 10, 11, 12]. In the parameter server architecture, a central server aggregates gradients from NN nodes and broadcasts the averaged update. Federated learning (FL) adapts this worker-server template under privacy and availability constraints, which exacerbate client drift and statistical heterogeneity [3] [13] [14]. Techniques such as client subsampling, proximal regularization, and adaptive aggregation have been proposed to mitigate these issues [5] [15] [16]. By contrast, synchronous data-parallel training in data centers employs collective communication (e.g., All-Reduce) to compute global averages without a central server, enabling the training of frontier large-scale models [17, 18, 19, 20]. Our work focuses on the communication bottleneck common to both centralized and federated regimes.
@@ IV All-Reduce Compatible TopKK
At iteration $t$, all NN nodes synchronize a random seed $s$, which deterministically generates a shared Gaussian projection matrix $\bm{V}\in\mathbb{R}^{n\times r},\text{with }\mathrm{vec}(\bm{V})\sim\mathcal{N}(\bm{0},\bm{I}_{nr})$. Next, a row-wise sketch of the global gradient can be achieved through:
@@
Communication Cost of ARC-Top-$K$. We analyze the communication overhead of row-sparsifying updates for an $m\times n$ parameter block replicated across NN nodes. At iteration $t$, each node selects $K$ rows, yielding sparsity $\mu\coloneqq K/m$, and transmits compressed updates through a collective primitive. In ARC-Top-$K$, a compact rank-$r$ sketch is additionally exchanged to align the selected row sets across nodes. “Communication per iteration” refers to the number of scalar entries transmitted per node per iteration. As summarized in Table I, the dense baseline requires $2mn$ entries using All-Reduce. Independent Top-$K$ instead relies on All-Gather to merge both values and indices, which incurs (N1)(nK+K)(N-1)(nK+K) entries and fails to preserve global contractivity. By contrast, Rand-$K$ achieves global contractivity in expectation while communicating $2Kn$ entries via All-Reduce. Finally, ARC-Top-$K$ restores global contractivity and maintains index-free All-Reduce through the shared sketch, with total communication volume $2Kn+2mr$, which further reduces to $2Kn+2m$ when $r=1$.
@@ VI-A Pre-Training on CIFAR
We evaluate ARC-Top-$K$ by training ResNet on CIFAR-10[35]. Each experiment is repeated 3 times with different random seeds. We report mean accuracy and standard deviation. ResNet-18 and ResNet-50 are trained for 200 epochs with a learning rate of $1\times 10^{-3}$, weight decay of $5\text{\hskip-1.42262pt}\times\text{\hskip-1.42262pt}10^{-4}$ and local batch size of 1616. All algorithms use sparsity $\mu\text{\hskip-1.42262pt}=\text{\hskip-1.42262pt}0.2$. For ARC-Top-$K$, we set projection dimension $r=4$. Table II reports the results, with Dense denoting the uncompressed baseline. The highest accuracy in each model–optimizer pair is highlighted in bold. Results are shown both without EF and with EF enabled.
@@ VI-D Further Study
Saving in communication bits. We quantify communication bits by pre-training LLaMA-130M for 10,00010{,}000 iterations with a global batch size of 256256; all remaining hyperparameters follow Section VI-C. Figure 2 plots training loss against cumulative communicated bits.In the figure, ARC-Top-$K$ reaches the target loss with fewer bits than the dense baseline. Furthermore, although ARC-Top-$K$ involves slightly more communication per iteration than Rand-$K$, its faster convergence yields lowest total communication to attain the same loss.

### v3

@@ I Introduction
The rise of large-scale machine learning has established distributed training as a fundamental paradigm in modern AI systems. In centralized cloud environments, massive datasets are partitioned and sharded across $N$ nodes (e.g., GPUs) within a data center to enable data-parallel training [1] [2]. In contrast, federated learning keeps data on $N$ client devices (e.g., smartphones) for privacy and regulatory compliance, while a coordinating server aggregates model updates instead of raw data [3, 4, 5]. Both paradigms give rise to the following distributed stochastic optimization problem:
@@
Here, $N$ denotes the number of computing nodes, $\bm{\xi}$ represents a random data sample drawn from the local distribution $\mathcal{D}_{i}$, and $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the local loss function at node $i$, which may be non-convex. While each node possesses its local data and loss function, they collaborate to minimize the global loss function $f(\bm{x})$ defined in problem (1).
In such distributed settings, each node $i$ can only access its local data and calculate local gradient $\nabla F_{i}(\bm{x};\bm{\xi})$ of its loss function during optimization. However, communication is required to obtain information from other nodes, which frequently becomes the primary bottleneck and dominates end-to-end training time. The associated cost scales with both the model dimension $d$ and the number of nodes $N$, rendering it prohibitive for large-scale learning tasks. Consequently, minimizing the volume of communicated bits per round is a critical problem for efficient and scalable distributed learning.
@@ I-B Main Results
Table I compares Rand-$K$, Top-$K$, and ARC-Top-$K$. The results show that ARC-Top-$K$ enables communication-efficient algorithms to achieve the same theoretical convergence rate and communication overhead as Rand-$K$, while matching the empirical performance of Top-$K$. All reported convergence rates are obtained in combination with EF21M [8], and the transient complexity reflects the number of iterations required for the asymptotic rate term to dominate. In the table, $N$ is the number of nodes, $T$ is the number of iterations, $\sigma$ is the standard deviation of the stochastic gradient noise. Tensors in all algorithm are compressed in 2-dimensional view of $m\times n$ while $K$ represents the number of selected rows in each tensor.
@@ II Related Works
Distributed Learning. Distributed optimization underpins large-scale training in both data-center clusters and cross-device networks [1] [2] [9, 10, 11, 12]. In the parameter server architecture, a central server aggregates gradients from $N$ nodes and broadcasts the averaged update. Federated learning (FL) adapts this worker-server template under privacy and availability constraints, which exacerbate client drift and statistical heterogeneity [3] [13] [14]. Techniques such as client subsampling, proximal regularization, and adaptive aggregation have been proposed to mitigate these issues [5] [15] [16]. By contrast, synchronous data-parallel training in data centers employs collective communication (e.g., All-Reduce) to compute global averages without a central server, enabling the training of frontier large-scale models [17, 18, 19, 20]. Our work focuses on the communication bottleneck common to both centralized and federated regimes.
@@ IV All-Reduce Compatible TopKK
At iteration $t$, all $N$ nodes synchronize a random seed $s$, which deterministically generates a shared Gaussian projection matrix $\bm{V}\in\mathbb{R}^{n\times r},\text{with }\mathrm{vec}(\bm{V})\sim\mathcal{N}(\bm{0},\bm{I}_{nr})$. Next, a row-wise sketch of the global gradient can be achieved through:
@@
Communication Cost of ARC-Top-$K$. We analyze the communication overhead of row-sparsifying updates for an $m\times n$ parameter block replicated across $N$ nodes. At iteration $t$, each node selects $K$ rows, yielding sparsity $\mu\coloneqq K/m$, and transmits compressed updates through a collective primitive. In ARC-Top-$K$, a compact rank-$r$ sketch is additionally exchanged to align the selected row sets across nodes. “Communication per iteration” refers to the number of scalar entries transmitted per node per iteration. As summarized in Table I, the dense baseline requires $2mn$ entries using All-Reduce. Independent Top-$K$ instead relies on All-Gather to merge both values and indices, which incurs (N1)(nK+K)(N-1)(nK+K) entries and fails to preserve global contractivity. By contrast, Rand-$K$ achieves global contractivity in expectation while communicating $2Kn$ entries via All-Reduce. Finally, ARC-Top-$K$ restores global contractivity and maintains index-free All-Reduce through the shared sketch, with total communication volume $2Kn+2mr$, which further reduces to $2Kn+2m$ when $r=1$.
@@ VI-A Pre-Training on CIFAR
We evaluate ARC-Top-$K$ by training ResNet on CIFAR-10[35]. Each experiment is repeated 3 times with different random seeds. We report mean accuracy and standard deviation. ResNet-18 and ResNet-50 are trained for 200 epochs with a learning rate of $1\times 10^{-3}$, weight decay of $5\text{\hskip-1.42262pt}\times\text{\hskip-1.42262pt}10^{-4}$ and local batch size of $16$. All algorithms use sparsity $\mu\text{\hskip-1.42262pt}=\text{\hskip-1.42262pt}0.2$. For ARC-Top-$K$, we set projection dimension $r=4$. Table II reports the results, with Dense denoting the uncompressed baseline. The highest accuracy in each model–optimizer pair is highlighted in bold. Results are shown both without EF and with EF enabled.
@@ VI-D Further Study
Saving in communication bits. We quantify communication bits by pre-training LLaMA-130M for 10,00010{,}000 iterations with a global batch size of $256$; all remaining hyperparameters follow Section VI-C. Figure 2 plots training loss against cumulative communicated bits.In the figure, ARC-Top-$K$ reaches the target loss with fewer bits than the dense baseline. Furthermore, although ARC-Top-$K$ involves slightly more communication per iteration than Rand-$K$, its faster convergence yields lowest total communication to attain the same loss.